# Exponentiations of ultrafilters

### v4

@@ 1   Introduction
A large class of problems in arithmetic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\ldots\}$.
@@
Analogously, we say that the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ defined by
@@
is PR if for all finite partitions $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$ there are $i\leq k$ and $x_{1},\ldots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\ldots,x_{n}\right)=0$.
In [8], Schur proved one of the first and most important results in Arithmetic Ramsey Theory.
@@ 2   Exponentiations of ultrafilters
Sums and products of ultrafilters will be denoted by $\oplus$,$\odot$ , ⊙ respectively. Those notion are related to that of tensor product of ultrafilters in $\beta\mathbb{N}$, which is defined as follows:
@@
Moreover, we will need to talk about extensions of functions to $\beta\mathbb{N}$: for $f:\mathbb{N}^{n}\rightarrow\mathbb{N}$, we will denote by $\widehat{f}$ its unique continuous extension $\widehat{f}:\beta(\mathbb{N}^{n})\rightarrow\beta\mathbb{N}$. For example, when $S(x,y)=x+y$ is the sum, we have that $\hat{S}(p\odot q)=p\oplus q$ for all $p,q\in\beta\mathbb{N}$. As usual, to simplify notations we identify each natural number $n\in\mathbb{N}$ with the principal idempotent it generates, so that $\mathbb{N}\subseteq\beta\mathbb{N}$.
As well-known, ultrafilters are often a fundamental tool to study PR problems. For configurations, the connection between ultrafilters and partition regularity reads as follows:
@@
The above fact leads to the following definition:
@@
Let $p\in\beta\mathbb{N}$ and let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$. We say that $p$ witnesses the partition regularity of the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ if for all $A\in p$ there exist $x_{1},\ldots,x_{n}\in A$ such that $f\left(x_{1},\ldots,x_{n}\right)=0$. In this case, we write $p\models\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$.
In the next sections we will use a few specific results from other papers. For the readers easiness, we recall them here. The first states that when the PR of different systems of equations are witnessed by the same ultrafilter, these configurations can be mixed.
@@
[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\ldots,x_{n}\right),g\left(y_{1},\ldots,y_{m}\right)$ be two functions such that $p\models\left(f_{1}\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)$. Then $p\models\left(f_{1}\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
@@
The third result we need is [6, Theorem A], that states that certain ultrafilter equations admits no solution.
@@
Finally, we will the non-PR of combined sums and exponentiations stated in [7, Theorem 6].
@@
We will provide some examples of nonprincipal ultrafilters $p,q$ satisfying $E_{1}(p,q)=E_{2}(q,p)$ after Proposition 2.10. To prove this proposition, it is helpful to use fractional powers of ultrafilters; to make proper sense of these powers, we prefer to state and prove Proposition 2.10 for ultrafilters in $\beta\mathbb{R}^{+}_{d}$, namely elements of the Stone-Čech compactification of +:=$\mathbb{R}^{+}:=\{x\in\mathbb{R}\mid x>0\}$ endowed with the discrete topology. We extend the exponentiations to $\beta\mathbb{R}_{d}$ as follows: given $p,q\in\beta\mathbb{R}^{+}_{d}$, we let:
@@
    $a^{p}=b^{q}\Leftrightarrow\left(\log_{a}(b)\right)p=q$;
@@
$(2)$ $\Rightarrow$)⇒ ) Consider the function $\log_{a}:\mathbb{R}^{+}\rightarrow\mathbb{R}$. A straightforward computation shows that $\widehat{\log_{a}}\left(a^{p}\right)=p$, hence our hypothesis entails that $p=\widehat{\log_{a}}\left(b^{q}\right)$. Now we observe that, for any $A\subseteq\mathbb{R}^{+}$,
$A\in\widehat{\log_{a}}\left(b^{q}\right)\Leftrightarrow(\log_{a})^{-1}(A)\in\left(b^{q}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in(\log_{a})^{-1}(A)\}\in q\}\in b\Leftrightarrow$
$\{y\in\mathbb{R}^{+}\mid b^{y}\in(\log_{a})^{-1}(A)\}\in q\Leftrightarrow\{y\in\mathbb{R}^{+}\mid\log_{a}\left(b^{y}\right)\in A\}\in q\Leftrightarrow$
$\{y\in\mathbb{R}^{+}\mid y\log_{a}(b)\in A\}\in q\Leftrightarrow A\in\left(\log_{a}(b)\right)q.$
Hence $p=\left(\log_{a}(b)\right)q$, as desired.
$\Leftarrow$)⇐ ) Just observe that $b^{q}=b^{\left(\log_{a}(b)\right)p}$, and $b^{\left(\log_{a}(b)\right)p}=E_{1}\left(b^{\left(\log_{a}(b)\right)},p\right)$ by $(1)$. We conclude as $b^{\left(\log_{a}(b)\right)}=a$.
$(3)$ $\Rightarrow$)⇒ ) This time, consider the function $f(x):=\sqrt[a]{x}:\mathbb{R}^{+}\rightarrow\mathbb{R}$. Again, a straightforward computation shows that $\widehat{f}\left(p^{a}\right)=p$, hence $p=\widehat{f}\left(q^{b}\right)$. For any $A\subseteq\mathbb{R}^{+}$ we have that
$A\in\widehat{f}\left(q^{b}\right)\Leftrightarrow(f)^{-1}(A)\in\left(q^{b}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid y^{x}\in f^{-1}(A)\}\in q\}\in b\Leftrightarrow$
$\{y\in\mathbb{R}^{+}\mid y^{b}\in f^{-1}(A)\}\in q\Leftrightarrow\{x\in\mathbb{R}^{+}\mid f\left(x^{b}\right)\in A\}\in q\Leftrightarrow$
$\{x\in\mathbb{R}^{+}\mid x^{\frac{b}{a}}\in A\}\in q\Leftrightarrow A\in q^{\frac{b}{a}}.$
$\Leftarrow$)⇐ ) By $(1)$ we have that $q^{b}=E_{1}\left(p^{\frac{a}{b}},b\right)=E_{1}\left(p,\frac{a}{b}\cdot b\right)=p^{a}$.
@@
Let $p,q,\beta\mathbb{N}\setminus{1}$, $a,b\in\mathbb{N}\setminus\{1\}$. Then
@@
Let $n\in\mathbb{N},n>1$ and $p,q\in\beta\mathbb{N}\setminus\{0\}$ be such that $p\odot q=q\odot p$. Then by Proposition 2.10.we have that
@@
In particular, this shows the existence of infinitely many pairs of nonprincipal ultrafilters solving $E_{1}(p,q)=E_{2}(q,p)$ (just take, e.g., $p=2^{q}$).
@@
Let $p\in\beta\mathbb{N}\setminus\{\mathbb{N}\}$, and let $a,b\in\mathbb{N}$. Then $p^{a}\odot p^{b}\neq p^{a+b}$.
@@
Let $F:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps a number $n$ to its largest prime divisor. Observe that, for all $p_{1},p_{2}\in\beta\mathbb{N}$, $\hat{\Omega}(E_{2}(p_{1},p_{2}))=p_{1}\odot\hat{\Omega}(p_{2})$ and $\hat{\Omega}(p_{1}\odot p_{2})=\hat{\Omega}(p_{1})\oplus\hat{\Omega}(p_{2})$. We consider two cases.
Case 1: $\hat{F}(p)=i\in\mathbb{N}$. In this case, consider the function $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ that maps a number $n$ to the sum of the exponents in its prime factorization. Necessarily, $\hat{\Omega}(p)\notin\mathbb{N}$; in fact, if by contrast $\hat{\Omega}(p)=j\in\mathbb{N}$, we would have that $p\in\overline{S}$, where
@@
which is a contradiction as $p$ is nonprincipal and $S$ is finite. Hence $\hat{\Omega}(p)\notin\mathbb{N}$. As $p^{a}\odot p^{b}=p^{a+b}$, we have that
$\hat{\Omega}\left(p^{a}\odot p^{b}\right)=\hat{\Omega}\left(p^{a}\right)\oplus\hat{\Omega}\left(p^{b}\right)=a\hat{\Omega}(p)\oplus b\hat{\Omega},$
@@
$\hat{\Omega}\left(p^{a+b}\right)=(a+b)\hat{\Omega}(p).$
We conclude as $a\hat{\Omega}(p)\oplus b\hat{\Omega}(p)\neq(a+b)\hat{\Omega}(p)$ by Theorem 2.6.
Case 2: $\hat{F}(p)\notin\mathbb{N}$. Let $G:\mathbb{N}\rightarrow\mathbb{N}$ be the function $n\rightarrow\max\{m\in\mathbb{N}\mid F(n)^{m}\ \text{divides}\ n\}$, and let $H(n):=F(n)^{G(n)}$. As $\hat{F}(p)$ is nonprincipal, also $\hat{H}(p)$ is nonprincipal. And
$\hat{H}\left(p^{a}\odot p^{b}\right)=b\hat{H}(p)\neq(a+b)\hat{H}(p)=\hat{H}\left(p^{a+b}\right),$
@@
The above results prove that the equalities $E_{1}(r,p)\odot E_{1}(r,q)=E_{1}(r,p\oplus q),E_{2}(p,r)\odot E_{2}(q,r)=E_{2}(p\oplus q,r)$ does not hold when $r$ is principal. They do not need to hold even when $r$ is nonprincipal. E.g., assume that $r$ is nonprincipal and $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}\in r$. A direct computation shows that
$I_{1}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}<a<p_{2}<b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{1}(r,p)\odot E_{1}(r,q),$
$I_{2}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}>a>p_{2}>b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{2}(p,r)\odot E_{2}(q,r),$
@@
$J=\{n\in\mathbb{N}\mid\exists p\in P,a\in\mathbb{N}\ \text{s.t.}\ p^{a}=n\}\in E_{1}(r,p\oplus q)\cap E_{2}(p\oplus q,r),$
and we conclude that $E_{1}(r,p)\odot E_{1}(r,q)\neq E_{1}(r,p\oplus q)$ and $E_{2}(p,r)\odot E_{2}(q,r)\neq E_{2}(p\oplus q,r)$ as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
Their definition should make evident that $E_{1},E_{2}$ are ultrafilters operations that can be used to study the partition regularity of certain exponential configurations.
Example 2.15.
Let $p\in\beta\mathbb{N}\setminus\mathbb{N}$ and let $q=E_{1}(p,p)$. Let $A\in q$. Then
$B_{0}=\{m\in\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m^{n}\in A\}\in p\}\in p.$
For $F_{1}\in\mathcal{P}_{f}(B_{0})$, let
$B_{1}=\bigcap_{t\in F_{1}}\{n\in\mathbb{N}\mid t^{n}\in A\}\cap B_{0}\in p.$
Proceeding inductively like above, it is possible to construct a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ such that, for all $j<i$, $F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A$. For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j<i\}\subseteq A$. Arguing as above with $E_{2}$ instead of $E_{1}$, we could deduce an analogous results with the order relationship between $i$ and $j$ switched.
The reader with a basic knowledge of ultrafilters will recognize that a trivial way to strengthen the above examples would be to take $p$ such that $E_{1}(p,p)=p$ or $E_{2}(p,p)=p$, namely some kind of exponential idempotent. We devote Section 2.3 to the study of the existence of such ultrafilters.
We now want to show that the operations $E_{1},E_{2}$ are related to the partition regularity of exponential triples, as well as of other exponential configurations. This should not come as a surprise: the main results in [2] can be restated in terminology introduced in Section 2.1 as follows:
@@
If $p$ is Van der Waerden’s witness then $2^{2^{p}\odot p}\models\{x,y,x^{y}\}$.
@@
  • if $p$ is also combinatorially rich, then $p\models x=y+z^{2}$ (see e.g. [1, Theorem 2.11], hence $q\models x=yt^{\log_{2}(t)}$,
@@
Let $q\in\beta\mathbb{N}\setminus\{0,1\}$. If $p$ is a Van der Waerden’s witness then $E_{1}\left(q,E_{1}(q,p)\odot p\right)\in\mathrm{ESW}$.
@@
Let $p,q\in\beta\mathbb{N}\setminus\{0,1\}$. Assume that, for all $n>1$, $p\models\{x,y,xn^{y}\}$. Then $E_{1}\left(q,p\right)\in\mathrm{ESW}$.
@@
Let $q\in\overline{K(\beta\mathbb{N},\odot)}$. Then for all $p\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $E_{1}\left(p,q\right)\in\mathrm{ESW}$.
@@
If $Sc_{+}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, Theorem 2.7 tells us that $Sc_{+}\cap\mathrm{ESW}=\emptyset$; in particular, if $E(\oplus)=\{p\in\beta\mathbb{N}\mid p\oplus p=p\}$ then
@@
as $\overline{K(\beta\mathbb{N},\odot)}\cup\overline{E(\oplus)}\subseteq Sc_{+}$. By Theorem 2.21.(1), we get that if $q\in\overline{K(\beta\mathbb{N},\odot)}$ then $E_{1}\left(p,q\right)\notin\overline{K(\beta\mathbb{N},\odot)}$; in particular, for $p,q\in\overline{K(\beta\mathbb{N},\odot)}$, $E_{1}(p,q)\notin\overline{K(\beta\mathbb{N},\odot)}$: the smallest bilateral multiplicative ideal of $\beta\mathbb{N}$ is very much not closed under the $E_{1}$ exponentiation.

### v5

@@ 1   Introduction
A large class of problems in arithmetic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\ldots\}$. We let $\omega=\mathbb{N}\cup\{0\}$.
@@
Analogously, if $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\omega$ and the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ is defined by
@@
we say that $\sigma\left(x_{1},\ldots,x_{n}\right)$ is PR if for every finite partition $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$, there are $i\leq k$ and $x_{1},\ldots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$.
In [8], Schur proved one of the first and most important results in rithmetic Ramsey theory.
@@ 2   Exponentiations of ultrafilters
Sums and products of ultrafilters will be denoted by $\oplus$,$\odot$ , ⊙ respectively. They are related to tensor products of ultrafilters which, over $\mathbb{N}$, are defined as follows:
@@
Moreover, we will need to talk about extensions of functions to $\beta\mathbb{N}$: for $f:\mathbb{N}^{n}\rightarrow\mathbb{N}$, we will denote by $\widehat{f}$ its unique continuous extension $\widehat{f}:\beta(\mathbb{N}^{n})\rightarrow\beta\mathbb{N}$. As usual, to simplify notations we identify each natural number $n\in\mathbb{N}$ with the principal idempotent it generates, so that $\mathbb{N}\subseteq\beta\mathbb{N}$, and we convene that, for $a\in\mathbb{N},p\in\beta\mathbb{N}$, we will write $ap$ instead of $a\odot p$. As in Proposition 2.13 we will use ultrafilters over $\mathbb{R}^{+}_{d}$, we recall that $\beta\mathbb{R}^{+}_{d}$ is the Stone-Čech compactification of +:=$\mathbb{R}^{+}:=\{x\in\mathbb{R}\mid x>0\}$ endowed with the discrete topology. We adopt for $\beta\mathbb{R}^{+}_{d}$ the analogous notational conventions we are using for $\beta\mathbb{N}$.
As well-known, ultrafilters are often a fundamental tool to study PR problems. For configurations, the connection between ultrafilters and partition regularity reads as follows (see e.g. [5, Theorem 5.7]):
@@
Analogously, for systems the following holds:
Fact 2.3.
Let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\omega$, and let the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ be defined by
$\begin{cases}f_{1}\left(x_{1},\ldots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\ldots,x_{n}\right)=0.\end{cases}$
The following are equivalent:
  1. 1.
    $\sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$ is PR;
  2. 2.
    $\exists p\in\beta\mathbb{N}\ \forall A\in p\ \exists x_{1},\ldots,x_{n}\in A\ \sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$.
The above facts lead to the following definitions:
@@
Let $p\in\beta\mathbb{N}$ and let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$. We say that $p$ witnesses the partition regularity of the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ if for all $A\in p$ there exist $x_{1},\ldots,x_{n}\in\mathbb{N}$ such that $f\left(x_{1},\ldots,x_{n}\right)\in A$. In this case, we write $p\models\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$. Analogously, if the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ is defined by
we say that $p$ witnesses the partition regularity of $\sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$ if $\forall A\in p\ \exists x_{1},\ldots,x_{n}\in A\ \sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$. In this case, we write $p\models\sigma\left(x_{1},\dots,x_{n}\right)=\vec{0}$.
In the next sections we will use a few specific results from other papers. For the readers’ easiness, we recall them here. The first states that when the PR of different systems of equations are witnessed by the same ultrafilter, these configurations can be mixed.
@@
[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\ldots,x_{n}\right),g\left(y_{1},\ldots,y_{m}\right)$ be two functions such that $p\models\left(f\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)$. Then $p\models\left(f\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
@@
The third and fourth result we need are [6, Theorem A] and (the restriction to $\mathbb{N}$ of) [5, Theorem 13.14] [^4] That we slightly rephrase to adapt to our notations., that states that certain ultrafilter equations admit no solution.
@@
Theorem 2.8.
[5, Theorem 13.14] Let $p,q,r,s\in\beta\mathbb{N}\setminus\mathbb{N}$. If |$|\{a\in\mathbb{N}\mid a\mathbb{Z}\in r\}|=|\mathbb{N}|$, then $q\oplus p\neq s\odot r$.
Finally, we will use the non-PR of combined sums and exponentiations stated in [7, Theorem 6].
@@
We will provide some examples of nonprincipal ultrafilters $p,q$ satisfying $E_{1}(p,q)=E_{2}(q,p)$ after Proposition 2.13. To prove this proposition, it is helpful to use fractional powers of ultrafilters; to make proper sense of these powers, we prefer to state and prove Proposition 2.13 for ultrafilters in $\beta\mathbb{R}^{+}_{d}$. We extend the exponentiations to $\beta\mathbb{R}_{d}$ as follows: given $p,q\in\beta\mathbb{R}^{+}_{d}$, we let:
@@
We start with a preliminary lemma generalizing standard laws of exponents and logarithms.
Lemma 2.12.
Let $p\in\beta\mathbb{R}_{d}^{+}$ and let $a,b\in\mathbb{R}^{+}\setminus\{1\}$. The following facts hold:
  1. 1.
    $\widehat{\log_{a}}\left(b^{q}\right)=\log_{a}(b)\odot q$;
  2. 2.
    $\left(a^{b}\right)^{q}=a^{bq}$;
  3. 3.
    $\left(p^{a}\right)^{b}=p^{ab}$.
Proof.
We prove only $(1)$, as $(2)-(3)$ are similar. Let $A\subseteq\mathbb{R}^{+}$. Then
$A\in\widehat{\log_{a}}\left(b^{q}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\log_{a}(x)\in A\}\in b^{q}\Leftrightarrow$
$\{z\in\mathbb{R}^{+}\mid b^{z}\in\{x\in\mathbb{R}^{+}\mid\log_{a}(x)\in A\}\}\in q\Leftrightarrow\{z\in\mathbb{R}^{+}\mid\log_{a}(b^{z})\in A\}\in q\Leftrightarrow$
$\{z\in\mathbb{R}^{+}\mid z\cdot\log_{a}(b)\in A\}\in q\Leftrightarrow A\in\log_{a}(b)\odot q.\qed$
@@
    $a^{p}=b^{q}\Leftrightarrow\left(\log_{a}(b)\right)\odot p=q$;
@@
$(2)$ If $a^{p}=b^{q}$, by Lemma 2.12.(1) $p=\widehat{\log_{a}}\left(a^{p}\right)=\widehat{\log_{a}}\left(b^{q}\right)=\left(\log_{a}(b)\right)\odot q$; conversely, if $p=\left(\log_{a}(b)\right)\odot q$ then $a^{p}=a^{\left(\log_{a}(b)\right)\odot q}=\left(a^{\log_{a}(b)}\right)^{q}=b^{q}$ by Lemma 2.12.(2).
$(3)$ If $p^{a}=q^{b}$, by Lemma 2.12.(3) $p^{\frac{a}{b}}=\left(p^{a}\right)^{\frac{1}{b}}=\left(q^{b}\right)^{\frac{1}{b}}=q$; conversely, again by Lemma 2.12.(3), if $p^{\frac{a}{b}}=q$ then $q^{b}=\left(p^{\frac{a}{b}}\right)^{b}=p^{\frac{ab}{b}}=p^{a}$.
@@
Let $p,q\in\beta\mathbb{N}\setminus{1}$, $a,b\in\mathbb{N}\setminus\{1\}$. Then
@@
Let $n\in\mathbb{N}\setminus\{1\}$ and $p,q\in\beta\mathbb{N}$ be such that $p\odot q=q\odot p$. Then by Proposition 2.13 we have that
@@
Let us notice that the equation $p\odot q=q\odot p$ can be solved by many pairs $p,q\in\beta\mathbb{N}\setminus\mathbb{N}$, for example by letting $p=q$, or by taking $p,q$ multiplicative idempotents with $p<q$. Hence, the equation $E_{1}(u,v)=E_{2}(u,v)$ has plenty of nontrivial solutions in $\beta\mathbb{N}\times\beta\mathbb{N}$.
@@
Let $p\in\beta\mathbb{N}\setminus\mathbb{N}$, and let $a,b\in\mathbb{N}$. Then $p^{a}\odot p^{b}\neq p^{a+b}$.
@@
Let $F:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps a number $n$ to its largest prime divisor. Observe that, for all $p_{1},p_{2}\in\beta\mathbb{N}$, $\widehat{F}(E_{2}(p_{1},p_{2}))=p_{1}\odot\widehat{F}(p_{2})$. We consider two cases.
Case 1: $\widehat{F}(p)=i\in\mathbb{N}$. In this case, consider the function $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ that maps a number $n$ to the sum of the exponents in its prime factorization. Necessarily, $\widehat{\Omega}(p)\notin\mathbb{N}$; in fact, if by contrast $\widehat{\Omega}(p)=j\in\mathbb{N}$, we would have that $p\in\overline{S}$, where
@@
which is a contradiction as $p$ is nonprincipal and $S$ is finite. Hence $\widehat{\Omega}(p)\notin\mathbb{N}$. As $p^{a}\odot p^{b}=p^{a+b}$, we have that
$\widehat{\Omega}\left(p^{a}\odot p^{b}\right)=\widehat{\Omega}\left(p^{a}\right)\oplus\widehat{\Omega}\left(p^{b}\right)=a\widehat{\Omega}(p)\oplus b\widehat{\Omega}(p),$
@@
$\widehat{\Omega}\left(p^{a+b}\right)=(a+b)\widehat{\Omega}(p).$
We conclude as $a\widehat{\Omega}(p)\oplus b\widehat{\Omega}(p)\neq(a+b)\widehat{\Omega}(p)$ by Theorem 2.7.
Case 2: $\widehat{F}(p)\notin\mathbb{N}$. Let $G:\mathbb{N}\rightarrow\mathbb{N}$ be the function $n\rightarrow\max\{m\in\mathbb{N}\mid F(n)^{m}\ \text{divides}\ n\}$, and let $H(n):=F(n)^{G(n)}$. As $\widehat{F}(p)$ is nonprincipal, also $\widehat{H}(p)$ is nonprincipal. Notice that:
  • $\widehat{H}\left(p^{a}\odot p^{b}\right)=b\widehat{H}(p)$: in fact, for $x\in\mathbb{N}$, let $D_{x}=\{y\in\mathbb{N}\mid F(y)>F(x)\}$. If $x\in\mathbb{N}$ and $q\in\overline{D_{x}}$, then $\widehat{H}(x\odot q)=\widehat{H}(q)$. Therefore, if we let $\rho_{q}:\beta\mathbb{N}\rightarrow\beta\mathbb{N}$ be the function mapping $u\in\beta\mathbb{N}$ into $u\odot q$, we have that $\widehat{H}\circ\rho_{q}$ is constantly equal to $\widehat{H}(q)$ on $\mathbb{N}$; as both these functions are continuous and $\mathbb{N}$ is dense in $\beta\mathbb{N}$, $\widehat{H}(p\odot q)=\widehat{H}(q)$. Therefore $\widehat{H}\left(p^{a}\odot p^{b}\right)=\widehat{H}\left(p^{b}\right)=b\widehat{H}(p)$;
  • $(a+b)\widehat{H}(p)=\widehat{H}\left(p^{a+b}\right)$, as for all $n\in\mathbb{N}$ $(a+b)H(n)=H\left(n^{a+b}\right)$.
Therefore
$\widehat{H}\left(p^{a}\odot p^{b}\right)=b\widehat{H}(p)\neq(a+b)\widehat{H}(p)=\widehat{H}\left(p^{a+b}\right),$
@@
The above result proves that the equalities $E_{1}(p,a)\odot E_{1}(p,b)=E_{1}(p,a+b),E_{2}(a,p)\odot E_{2}(b,p)=E_{2}(a+b,p)$ does not hold when $a,b\in\mathbb{N},p\in\beta\mathbb{N}\setminus\mathbb{N}$. They do not need to hold even when $a,b\in\beta\mathbb{N}\setminus\mathbb{N}$. E.g., assume that $p,q,r$ are nonprincipal and $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}\in p$. A direct computation shows that
$I_{1}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}<a<p_{2}<b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{1}(p,q)\odot E_{1}(p,r),$
$I_{2}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}>a>p_{2}>b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{2}(q,p)\odot E_{2}(r,p),$
@@
$J=\{n\in\mathbb{N}\mid\exists p\in P,a\in\mathbb{N}\ \text{s.t.}\ p^{a}=n\}\in E_{1}(p,q\oplus r)\cap E_{2}(q\oplus r,p),$
and we conclude that $E_{1}(p,q)\odot E_{1}(p,r)\neq E_{1}(p,q\oplus r)$ and $E_{2}(q,p)\odot E_{2}(r,p)\neq E_{2}(q\oplus r,p)$ as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
We now want to show that the operations $E_{1},E_{2}$ are related to the partition regularity of exponential triples, as well as of other exponential configurations. This should not come as a surprise: the main results in [2] can be restated in the terminology introduced in Section 2.1 as follows:
@@
If $p$ is a Van der Waerden’s witness then $2^{2^{p}\odot p}\models\{x,y,x^{y}\}$.
@@
  • if $p$ is also combinatorially rich [^6] See e.g. [1, Section 1.3] for a definition., then $p\models x=y+z^{2}$ (see e.g. [1, Theorem 2.11]), hence $q\models x=yt^{\log_{2}(t)}$,
@@
Let $q\in\beta\mathbb{N}\setminus\{1\}$. If $p$ is a Van der Waerden’s witness then $E_{1}\left(q,E_{1}(q,p)\odot p\right)\in\mathrm{ESW}$.
@@
Let $p,q\in\beta\mathbb{N}\setminus\{1\}$. Assume that, for all $n>1$, $p\models\{x,y,xn^{y}\}$. Then $E_{1}\left(q,p\right)\in\mathrm{ESW}$.
@@
Let $q\in\overline{K(\beta\mathbb{N},\odot)}$. Then for all $p\in\beta\mathbb{N}\setminus\{1\}$ we have that $E_{1}\left(p,q\right)\in\mathrm{ESW}$.
@@
If $\mathrm{ASW}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, Theorem 2.9 tells us that $\mathrm{ASW}\cap\mathrm{ESW}=\emptyset$; in particular, if $E(\oplus)=\{p\in\beta\mathbb{N}\mid p\oplus p=p\}$ then
@@
as $\overline{K(\beta\mathbb{N},\odot)}\cup\overline{E(\oplus)}\subseteq\mathrm{ASW}$. By Theorem 2.23.(1), we get that if $q\in\overline{K(\beta\mathbb{N},\odot)}$ then $E_{1}\left(p,q\right)\notin\overline{K(\beta\mathbb{N},\odot)}$; in particular, for $p,q\in\overline{K(\beta\mathbb{N},\odot)}$, $E_{1}(p,q)\notin\overline{K(\beta\mathbb{N},\odot)}$: the smallest bilateral multiplicative ideal of $\beta\mathbb{N}$ is very much not closed under the $E_{1}$ exponentiation.